# Performance in solving the Hermitian and pseudo-Hermitian Bethe-Salpeter equation with the Yambo code

### v2

@@ 3.1 Details on the Performed Simulations
The characteristics of two clusters used for CPU-only and GPU simulations are reported in Table 1. We constructed matrices from Eq. (1) in single precision (using 32-bit float numbers for both real and imaginary parts) for the case $A=R$ for a single layer of chromium triiodide CrI3. The set of BSE-like matrices considered in the present report are generated by replacing the demanding BSE kernel with the cheaper TD-DFT kernel [1, 43], as implemented in the Yambo code. Specifically, the matrix blocks are defined as follows:
@@
Here, $i=\{nm\mathbf{k}\}$ is a transition index. In the resonant block both $i$ and $i^{\prime}$ refer to positive energy transitions from occupied to empty states, $n\mathbf{k}\rightarrow m\mathbf{k}$, while in the coupling block $j^{\prime}$ refers to negative energy transitions, $m\mathbf{k}\rightarrow n\mathbf{k}$, and $\Delta\epsilon_{i}=\epsilon_{m\mathbf{k}}-\epsilon_{n\mathbf{k}}$ are energy differences. We used an energy cut-off of 4 Ry for the eh-exchange $v$, and 10 Ry for the $f_{xc}$ kernel. These parameters determine the prefactor for the building time of the BSE kernel which we do not analyze here. The size of the BSE matrix is defined by the number of bands used. We considered the three cases reported in Table 2. Here we focus on the performance of the two solvers described in Sect. 2. For direct diagonalization, ScaLAPACK and ELPA libraries (solver type 2 on CPU, type 1 on GPU) were used with a block size of $64\times 64$ for the block-cyclic distribution. For the iterative solver, the SLEPc library was used, with either explicit (e) or shell (s) matrix formats. In our experience, when the eigenvectors are extracted, if more than 5% of the spectrum is required, direct diagonalization routines are recommended. For smaller fractions of the spectrum, the SLEPc iterative solvers are preferable. Here we fix the number of extracted eigenvectors to 100. In the parallel runs this ranges from 1% (for the smallest matrices), to 0.1% (for the largest considered matrix) of the spectrum.
@@
For an efficiency plot, when a reference value is not available, we estimate it using the exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, where $t_{1}$ represents the time required for a run on a single MPI task, and $p$ is a constant that characterizes how the simulation time scales with the number of MPI tasks  [44].
For perfect scaling, $p=-1$, i.e., doubling the number of MPI tasks should halve the simulation time. Any value $p>-1$ indicates inefficiencies in parallelization, with a time reduction lower than expected. We will refer to the fitted parameter $p$ as the parallelization factor.
For both CPU and GPU computations, we measured memory usage using the Linux time command, which reports peak memory of each MPI task. For some of the runs we also monitored the peak memory of the run reported by the PETSc internal memory tracking tools. To analyze the memory use, we assume that the allocated host memory is due to arrays/matrices which are either perfectly distributed or fully duplicated. We define the distributed memory as the intercept, and the duplicated memory as the slope of the linear fit of the allocated memory vs number of MPI tasks. Moreover, since in some simulations we observe a sudden memory jump from the serial to the parallel simulations, we also define the parallel–overhead memory as the difference between the fit intercept and the memory allocated in the serial run. We attribute this additional memory increase in parallel runs to the allocation of extra variables not present during serial execution. Memory extracted by the time command and the PETSc tracker agree well, with a parallel overhead memory detected only by the time command. [45]
@@ 3.2 Comparison of non-Hermitian vs pseudo-Hermitian Algorithms
Figure 1: Time complexity (panels (a)–(c)) and memory complexity (panels (d)–(f)) for solving the BSE eigenvalue problem using a single CPU. Results are shown for: (a,d) the Hermitian case, (b,e) the coupling case with a non-Hermitian algorithm, and (c,f) the coupling case with a pseudo-Hermitian algorithm. In all cases, $N$ denotes the size of the resonant block. Simulations were carried out on the ISM cluster.
Before analyzing the CPU and GPU scaling of the simulations, we first analyze how the computational time scales with matrix size (i.e., time complexity) using a single CPU. We compare the standard algorithms for Hermitian (H) matrices (corresponding to resonant case), standard algorithms for non-Hermitian matrices (NH), and novel algorithms that exploit the pseudo-Hermitian (PH) structure of the coupling matrix. Here we consider three different solvers: (i) full diagonalization which is based on (Sca)LAPACK [46] (see Sect. 2.1.1), (ii) Krylov-Schur iterative solver based on the SLEPc library (see Sect. 2.2), and (iii) Haydock iterative solver. The Haydock solver performance is reported for reference as an example of an iterative algorithm implemented in Yambo, which already exploits the pseudo-Hermitian structure of the matrix [41].
Fig. 1 presents time complexity plots, and Table 3 reports the corresponding slope and intercept values, $log(t)=A\cdot log(N)+B$. By moving from a logarithmic representation to the direct time versus matrix size relation, the slope A corresponds to the scaling exponent of the algorithm with respect to matrix size. Lower values of A indicate better scaling. The intercept B represents the computational prefactor, which is also more favorable when smaller. The relation is expressed as $t=B\,N^{A}$. The expected scaling is $\mathcal{O}(N^{2})$ for SLEPc and $\mathcal{O}(N^{3})$ for exact diagonalization. Fig. 1 shows that full diagonalization is the slowest algorithm, while Haydock is the fastest, and SLEPc is intermediate. The performance differences between SLEPc and Haydock arise from the fact that the Haydock algorithm only extracts the optical spectrum, while the SLEPc algorithm also extracts a few eigenvalues and eigenvectors [11].
@@
Table 3: Slope (A) and intercept (B) parameters from linear regressions of the time-complexity data for the resonant (H) and coupling cases (non-Hermitian (NH) and pseudo-Hermitian (PH)). The upper section reports results obtained on the ISM cluster using the GFortran compiler; these are the data shown in Fig. 1. The lower section presents reproduced benchmarks performed on the University of Valencia cluster (UVC) using Intel compilers and MKL libraries.
By comparing panels (b) and (c) of Fig. 1, we observe a significant reduction in the time to solution when switching from the NH to the PH algorithm, for both exact diagonalization and the SLEPc based iterative solver. For the SLEPc case a five- to six-fold improvement in speed is obtained. As shown in Table 3, the improvements are due to the better scaling exponential (i.e., slope). In the NH implementation, the scaling behavior was similar to that of full diagonalization algorithms, and the performance advantage over diagonalization was mainly due to a smaller prefactor (i.e., a lower intercept in the scaling fit). For the diagonalization case the speed-up from the NH to the PH case is also roughly a factor five, while the improvement is due to both a slightly better scaling and prefactor [47].
@@
Panels (d), (e), and (f) in Fig. 1 and Table 4 illustrate the memory complexity of the studied algorithms, that is, how memory usage increases with matrix size. The following observations can be made: (i) iterative algorithms consume less memory than diagonalization-based algorithms; (ii) memory complexity is similar between LAPACK and ELPA, and between SLEPc and Haydock; and (iii) the PH implementations result in improved memory efficiency vs the NH implementation for the SLEPc algorithms, while causing a slight increase in memory usage for the LAPACK-based approach [48]. Panels (d) and (f) show that the ELPA and LAPACK memory fits overlap almost perfectly. In contrast, a small difference is observed between Haydock and SLEPc, with the Haydock algorithm being slightly more memory-efficient. Overall the memory increase as a function of the matrix size behaves as expected.
@@
Figure 2: Comparison of time complexity for the studied algorithms, with the x-axis representing the actual matrix size: N for the resonant case and 2N for the coupling case.
Fig. 2 shows the time complexity plots rescaled to reflect the actual size of the coupling matrix. As already observed, iterative solvers are orders faster than diagonalization-based approaches. When the full matrix size is considered, instead of only the resonant block size, it can be seen that the two algorithms (H vs PH) behave rather similar performances. The ELPA implementation is the only one where the PH is clearly faster than the H for all matrix sizes considered here. For the iterative solvers, the Hermitian version exhibits a more favorable slope compared to the pseudo-Hermitian one, and this difference becomes more significant as the matrix size increases.
@@
Figure 3: Execution time (panels (a) and (b)), efficiency (panels (c) and (d)), and memory usage (panels (e) and (f)) for different solvers applied to the resonant and coupling cases of the BSE eigenvalue problem. CPU case.
Now we analyze MPI scaling of the solvers on CPU. We consider strong scaling with matrices of size $N=10^{4}$ and $N=3\cdot 10^{4}$. For the coupling case, this corresponds to matrices of size $2N$. In Fig. 3 time and memory scaling plots are presented, together with efficiency plots. The exponential fit of the data is reported in Table 5. Finally, the extracted parameters from the memory fit are reported in Table 6. In this and next subsection, the term scaling is referring to scaling with respect to number of CPUs or GPUs.
@@
Table 5: Exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, of time vs number of CPUs for resonant (res.) and coupling (cpl.) matrices. Where $p=-1$ is the ideal parallelization case. (e) – explicit matrix type, (s) – shell matrix type.
@@
Analyzing in detail different flavours of the iterative and diagonalization solvers we observe the following. For the SLEPc case, the shell matrix algorithm performs very poorly, with a total time which starts to grow when a critical number of MPI tasks is reached. This behaviour is somewhat expected as the shell approach uses a matrix distribution which is optimized for building the kernel. Instead, the explicit approach is more efficient, although not as efficient as full diagonalization. For the diagonalization case, the ELPA library outperforms ScaLAPACK, likely due to its more efficient parallelization. This performance difference can be seen in Fig. 3, panel (a). As the matrix size increases, ELPA’s efficiency will plateau in the range of 0.6 – 0.8, depending on the matrix type and size, with larger matrices benefiting more from ELPA’s optimization, whereas for ScaLAPACK, the plateau the plateau consistently occurs around 0.4 – 0.5. In the coupling case, panel (b), ELPA shows an even greater performance advantage. This is most likely due to its more optimized implementation of the pseudo-Hermitian solver, compared to ScaLAPACK. Interestingly, in some cases, ELPA’s efficiency in the pseudo-Hermitian case is higher than in the Hermitian case. These differences are also reflected in the p values reported in Table 5, where ELPA values are closer to the ideal p = $-1$, typically around $-0.8$ to $-0.9$, followed by ScaLAPACK with values in between $-0.7$ and $-0.8$, and finally the SLEPc explicit solver reaches values in between $-0.6$ and $-0.7$. The SLEPc shell solver, again, shows the worst performance in this comparison.
@@
Figure 4: Execution time (panels (a) and (b)), efficiency (panels (c) and (d)), and memory usage (panels (e) and (f)) for different solvers applied to the resonant and coupling cases of the BSE eigenvalue problem. GPU case. Only host memory is reported here.
@@
Table 7: Exponential fit of time vs number of GPUs, $t_{i}=t_{1}\,n_{i}^{p}$. The case $p=-1$ represents the ideal parallelization. The results marked with may be influenced by the limited number of data points in the low GPU count range.
@@
Memory usage analysis, panels (e) and (f), and Table 8 shows that duplicated memory remains consistently around 2 GB per GPU, which is 4 times higher than for the CPU case. The distributed memory is very close to the one obtained in the CPU case, see Table 6. On Leonardo Booster, per each GPU there are 64 GB of memory, and for the cases of medium and big matrices, this imposes a minimal amount of GPUs needed to perform the simulations. In comparison with the CPU case, on GPUs the SLEPc solver shows similar or even higher memory consumption than the ELPA solver. This is due to the large parallel overhead memory that we are presently investigating. For the GPU implementation, the parallel overhead memory is of a similar order of magnitude to that observed in the CPU case. Specifically, for SLEPc, the parallel overhead memory is large and positive, indicating the allocation of additional variables for parallel runs. Conversely, for ELPA, the parallel overhead is negative, suggesting more efficient parallelization schemes.

### v3

@@ 3.1 Details on the Performed Simulations
Node interconnect ConnectX®-4 VPI Adapter Card, FDR IB (56Gb/s) 2x dual-port Infiniband HDR interconnect network interface (400 Gbps aggregated)
@@
The characteristics of two clusters used for CPU-only and GPU simulations are reported in Table 1. We constructed matrices from Eq. (1) in single precision (using 32-bit float numbers for both real and imaginary parts) for the case $A=R$ for a single layer of chromium triiodide CrI3. The set of BSE-like matrices considered in the present report are generated by replacing the demanding BSE kernel with the cheaper TD-DFT kernel [Onida2002, marques2006time], as implemented in the Yambo code. Specifically, the matrix blocks are defined as follows:
@@
Here, $i=\{nm\mathbf{k}\}$ is a transition index. In the resonant block both $i$ and $i^{\prime}$ refer to positive energy transitions from occupied to empty states, $n\mathbf{k}\rightarrow m\mathbf{k}$, while in the coupling block $j^{\prime}$ refers to negative energy transitions, $m\mathbf{k}\rightarrow n\mathbf{k}$, and $\Delta\epsilon_{i}=\epsilon_{m\mathbf{k}}-\epsilon_{n\mathbf{k}}$ are energy differences. We used an energy cut-off of 4 Ry for the eh-exchange $v$, and 10 Ry for the $f_{xc}$ kernel. These parameters determine the prefactor for the building time of the BSE kernel which we do not analyze here. The size of the BSE matrix is defined by the number of bands used. We considered the three cases reported in Table 2. Here we focus on the performance of the two solvers described in Sect. 2. For direct diagonalization, ScaLAPACK and ELPA libraries (solver type 2 on CPU, type 1 on GPU) were used with a block size of $64\times 64$ for the block-cyclic distribution. For the iterative solver, the SLEPc library was used, with either explicit (e) or shell (s) matrix formats. In our experience, when the eigenvectors are extracted, if more than 5% of the spectrum is required, direct diagonalization routines are recommended. For smaller fractions of the spectrum, the SLEPc iterative solvers are preferable. Here we fix the number of extracted eigenvectors to 100. In the parallel runs this ranges from 1% (for the smallest matrices), to 0.1% (for the largest considered matrix) of the spectrum. We selected a fixed number of 100 eigenvalues because, when using the SLEPc solver, the primary interest typically lies in the characterization of the lowest excitonic peak. In practice, this number is more than sufficient to capture the relevant low-energy excitations. If additional eigenvectors are required, for instance, to investigate higher excitonic peaks or a broader portion of the spectrum, then one must resort to full diagonalization solvers, which provide access to the complete set of excitonic states.
@@
For an efficiency plot, when a reference value is not available, we estimate it using the exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, where $t_{1}$ represents the time required for a run on a single MPI task, and $p$ is a constant that characterizes how the simulation time scales with the number of MPI tasks  [footnote4].
For perfect scaling, p=1p=-1, i.e., doubling the number of MPI tasks should halve the simulation time. Any value p>1p>-1 indicates inefficiencies in parallelization, with a time reduction lower than expected. We will refer to the fitted parameter $p$ as the parallelization factor.
For both CPU and GPU computations, we measured memory usage using the Linux time command, which reports peak memory on host for each MPI task. For some of the runs we also monitored the peak memory of the run reported by the PETSc internal memory tracking tools. To analyze the memory use, we assume that the allocated host memory is due to arrays/matrices which are either perfectly distributed or fully duplicated. We define the distributed memory as the intercept, and the duplicated memory as the slope of the linear fit of the allocated memory vs number of MPI tasks. Moreover, since in some simulations we observe a sudden memory jump from the serial to the parallel simulations, we also define the parallel–overhead memory as the difference between the fit intercept and the memory allocated in the serial run. We attribute this additional memory increase in parallel runs to the allocation of extra variables not present during serial execution. Memory extracted by the time command and the PETSc tracker agree well, with a parallel overhead memory detected only by the time command [footnote5].
The conclusions presented in this work are expected to be largely independent of the specific material. CrI3 is employed solely as a representative case for generating a dense BSE matrix. The only aspect that can vary from case to case is the convergence criteria. The full diagonalization algorithm scales as $N^{3}$ and the time to solution does not depend on the matrix properties. For the iterative SLEPc solver, the time required for a single iteration is also independent of the matrix properties, while the number of iterations needed might be affected by the matrix properties. For instance, if the matrix has clustered eigenvalues, this will hinder convergence, resulting in a larger number of iterations. However, this does not usually occur in BSE applications, and the chosen matrix gives a good representation of the needed time to solution for a wide range of cases. Apart from this, once the matrix is constructed, the diagonalization procedure itself is general and does not depend on the underlying material.
@@ 3.2 Comparison of non-Hermitian vs pseudo-Hermitian Algorithms
Figure 1: Time complexity (panels (a)–(c)) and memory complexity (panels (d)–(f)) for solving the BSE eigenvalue problem using a single CPU. Results are shown for: (a,d) the Hermitian case, (b,e) the coupling case with a non-Hermitian algorithm, and (c,f) the coupling case with a pseudo-Hermitian algorithm. In all cases, NN denotes the size of the resonant block. Simulations were carried out on the ISM cluster.
Before analyzing the CPU and GPU scaling of the simulations, we first analyze how the computational time scales with matrix size (i.e., time complexity) using a single CPU. We compare the standard algorithms for Hermitian (H) matrices (corresponding to resonant case), standard algorithms for non-Hermitian matrices (NH), and novel algorithms that exploit the pseudo-Hermitian (PH) structure of the coupling matrix. Here we consider three different solvers: (i) full diagonalization which is based on (Sca)LAPACK [footnote6] (see Sect. 2.1.1), (ii) Krylov-Schur iterative solver based on the SLEPc library (see Sect. 2.2), and (iii) Haydock iterative solver. The Haydock solver performance is reported for reference as an example of an iterative algorithm implemented in Yambo, which already exploits the pseudo-Hermitian structure of the matrix [Gruning:2011:ITL].
Fig. 1 presents time complexity plots, and Table 3 reports the corresponding slope and intercept values, $\log(t)=A\cdot\log(N)+B$. By moving from a logarithmic representation to the direct time versus matrix size relation, the slope A corresponds to the scaling exponent of the algorithm with respect to matrix size. Lower values of $A$ indicate better scaling. The intercept $B$ represents the computational prefactor, which is also more favorable when smaller. The relation is expressed as $t=B\,N^{A}$. The expected scaling is $\mathcal{O}(N^{2})$ for SLEPc and $\mathcal{O}(N^{3})$ for exact diagonalization. Fig. 1 shows that full diagonalization is the slowest algorithm, while Haydock is the fastest, and SLEPc is intermediate. The performance differences between SLEPc and Haydock arise from the fact that the Haydock algorithm only extracts the optical spectrum, while the SLEPc algorithm also extracts a few eigenvalues and eigenvectors [Sangalli2019].
Table 3: Slope (AA) and intercept ($B$) parameters from linear regressions of the time-complexity data for the resonant (H) and coupling cases (non-Hermitian (NH) and pseudo-Hermitian (PH)). The upper section reports results obtained on the ISM cluster using the GFortran compiler; these are the data shown in Fig. 1. The lower section presents reproduced benchmarks performed on the University of Valencia cluster (UVC) using Intel compilers and MKL libraries.
@@
By comparing panels (b) and (c) of Fig. 1, we observe a significant reduction in the time to solution when switching from the NH to the PH algorithm, for both exact diagonalization and the SLEPc based iterative solver. For the SLEPc case a five- to six-fold improvement in speed is obtained. As shown in Table 3, the improvements are due to the better scaling exponential (i.e., slope). In the NH implementation, the scaling behavior was similar to that of full diagonalization algorithms, and the performance advantage over diagonalization was mainly due to a smaller prefactor (i.e., a lower intercept in the scaling fit). For the diagonalization case the speed-up from the NH to the PH case is also roughly a factor five, while the improvement is due to both a slightly better scaling and prefactor [footnote7].
@@
Panels (d), (e), and (f) in Fig. 1 and Table 4 illustrate the memory complexity of the studied algorithms, that is, how memory usage increases with matrix size. The following observations can be made: (i) iterative algorithms consume less memory than diagonalization-based algorithms; (ii) memory complexity is similar between LAPACK and ELPA, and between SLEPc and Haydock; and (iii) the PH implementations result in improved memory efficiency vs the NH implementation for the SLEPc algorithms, while causing a slight increase in memory usage for the LAPACK-based approach [footnote8]. Panels (d) and (f) show that the ELPA and LAPACK memory fits overlap almost perfectly. In contrast, a small difference is observed between Haydock and SLEPc, with the Haydock algorithm being slightly more memory-efficient. Overall the memory increase as a function of the matrix size behaves as expected.
@@
Figure 2: Comparison of time complexity for the studied algorithms, with the x-axis representing the actual matrix size: NN for the resonant case and $2N$ for the coupling case.
Fig. 2 shows the time complexity plots rescaled to reflect the actual size of the coupling matrix. As already observed, iterative solvers are orders faster than diagonalization-based approaches. When the full matrix size is considered, instead of only the resonant block size, it can be seen that the two algorithms (H vs PH) have rather similar performances. The ELPA implementation is the only one where the PH is clearly faster than the H for all matrix sizes considered here. For the iterative solvers, the Hermitian version exhibits a more favorable slope compared to the pseudo-Hermitian one, and this difference becomes more significant as the matrix size increases.
@@
Figure 3: Execution time (panels (a) and (b)), efficiency (panels (c) and (d)), and memory usage (panels (e) and (f)) for different solvers applied to the resonant and coupling cases (pseudo-Hermitian solver) of the BSE eigenvalue problem. CPU case.
Now we analyze MPI scaling of the solvers on CPU. We consider strong scaling with matrices of size $N=10^{4}$ and $N=3\cdot 10^{4}$. For the coupling case, this corresponds to matrices of size $2N$. In Fig. 3 time and memory scaling plots are presented, together with efficiency plots. The exponential fit of the data is reported in Table 5. Finally, the extracted parameters from the memory fit are reported in Table 6. In this and next subsection, the term scaling is referring to scaling with respect to number of CPUs or GPUs. Moreover, only the pseudo-Hermitian version of the algorithm is considered for the coupling matrices.
Table 5: Exponential fit, $t_{i}=t_{1}\,n_{i}^{p}$, of time vs number of CPUs for resonant (res.) and coupling (cpl.) matrices. Where p=1p=-1 is the ideal parallelization case. (e) – explicit matrix type, (s) – shell matrix type.
@@
Analyzing in detail different flavours of the iterative and diagonalization solvers we observe the following. For the SLEPc case, the shell matrix algorithm performs very poorly, with a total time which starts to grow when a critical number of MPI tasks is reached. This behaviour is somewhat expected as the shell approach uses a matrix distribution which is optimized for building the kernel. Instead, the explicit approach is more efficient, although not as efficient as full diagonalization. For the diagonalization case, the ELPA library outperforms ScaLAPACK, likely due to its more efficient parallelization. This performance difference can be seen in Fig. 3, panel (a). As the matrix size increases, ELPA’s efficiency will plateau in the range of 0.6 – 0.8, depending on the matrix type and size, with larger matrices benefiting more from ELPA’s optimization, whereas for ScaLAPACK, the plateau the plateau consistently occurs around 0.4 – 0.5. In the coupling case, panel (b), ELPA shows an even greater performance advantage. This is most likely due to its more optimized implementation of the pseudo-Hermitian solver, compared to ScaLAPACK. Interestingly, in some cases, ELPA’s efficiency in the pseudo-Hermitian case is higher than in the Hermitian case. These differences are also reflected in the p values reported in Table 5, where ELPA values are closer to the ideal p = 1-1, typically around 0.8-0.8 to 0.9-0.9, followed by ScaLAPACK with values in between 0.7-0.7 and 0.8-0.8, and finally the SLEPc explicit solver reaches values in between 0.6-0.6 and 0.7-0.7. The SLEPc shell solver, again, shows the worst performance in this comparison.
@@
Figure 4: Execution time (panels (a) and (b)), efficiency (panels (c) and (d)), and memory usage (panels (e) and (f)) for different solvers applied to the resonant and coupling cases (pseudo-Hermitian solver) of the BSE eigenvalue problem. GPU case. Only host memory is reported here.
Table 7: Exponential fit of time vs number of GPUs, $t_{i}=t_{1}\,n_{i}^{p}$. The case p=1p=-1 represents the ideal parallelization. The results marked with may be influenced by the limited number of data points in the low GPU count range.
@@
Memory usage analysis, panels (e) and (f), and Table 8 shows that duplicated memory on the host remains consistently around 2 GB per used GPU, which is 4 times higher than for the CPU case. The distributed memory is very close to the one obtained in the CPU case, see Table 6. On Leonardo Booster, per each GPU there are 64 GB of memory, and for the cases of medium and big matrices, this imposes a minimal amount of GPUs needed to perform the simulations. In comparison with the CPU case, on GPUs the SLEPc solver shows similar or even higher memory consumption than the ELPA solver. This is due to the large parallel overhead memory that we are presently investigating. For the GPU implementation, the parallel overhead memory is of a similar order of magnitude to that observed in the CPU case. Specifically, for SLEPc, the parallel overhead memory is large and positive, indicating the allocation of additional variables for parallel runs. Conversely, for ELPA, the parallel overhead is negative, suggesting more efficient parallelization schemes.